# Identification and estimation of vaccine effectiveness in the test-negative design under equi-confounding

### v2

@@ Appendix D Doubly robust estimator under cohort sampling
$E\left[\Pr[Y=2|V=0,X]\dfrac{\Pr[Y=1|V=1,X]}{\Pr[Y=1|V=0,X]}\Big{|}V=1\right]$ (A8)
@@
$E\left[\dfrac{(1-V)}{\Pr(V=1)}\mathbbm{1}(Y=2)\dfrac{\pi(1,X)}{1-\pi(1,X)}\right]$ (A9)
@@ Appendix E Doubly robust estimator under TND sampling
Under TND sampling, i.e. $O_{TND}=\{(X_{i},V_{i},S_{i}=1,Y_{i}):i=1,\ldots,n\},$ with selection $S=\mathbbm{1}(Y\neq 0)$, we show in the main text that Equations A8 and A9, and by extension Equations A5 and A6, are still identifiable by
@@ G.2 Immortal time among vaccinated
Consider the emulation of a trial where events occurring in the first 28 days are excluded in both the vaccine arm and the control arm, as was specified in the original trials of the SARS-CoV-2 vaccines. This is often done because, biologically, it is believed that during this window the vaccine has not had time to provoke a sufficient immunological response and therefore provides no protection, although discarding these events may produce selection bias due to depletion of susceptibles when there is an effect. Regardless, attempts to emulate this result using a TND are challenged by the fact that the TND does not have a well-defined start of follow up and only the timing of vaccination and testing are known. Therefore, investigators instead discard cases where vaccination occurred less than 28 days prior to receiving a test, i.e. only among the vaccinated. This results in a form of immortal time bias as the vaccinated cannot have the outcome during this 28 day period. Assume for a second, to simpilify exposition, that conditional on $X$ there is no other source of confounding. Note that under our framework, this would imply that
@@ Appendix H Identification when test is imperfect
In the main text, we assumed the availability of a perfect test such that $\widetilde{Y}_{i}=Y_{i}$ for all individuals where $\widetilde{Y}$ is the test result and $Y$ is the true value. In the absence of such a test, here we describe the potential for bias due to misclassification of case status which has also been described previously [sullivan_theoretical_2016]. Because most studies employ real-time reverse-transcription polymerase chain reaction (RT-PCR), false positives are unlikely as most RT-PCR tests have specificity approaching 100%. Therefore, misclassification of test-positives is probably rare, perhaps due to sample contamination or data entry errors. False negatives may be be more likely as test sensitivity is generally lower due to the fact that (1) some of those infected with test-positive illness may not shed detectable virus or viral RNA; (2) some may seek care after shedding has ceased; or (3) sample quality may be compromised due to swab quality or inadequate storage. If sample collection is sufficient such that the source of the test-negative infection can be identified via RT-PCR, then sensitivity with respect to the test-positive illness can be improved by limiting to test-negative controls with an identified cause.
@@
    No effect of vaccination on test result That is, for any $v$ and $v^{\prime}$
@@ Appendix I Simulation details
where $\mu_{v}(X)=\Pr(Y=2|X,V=v)$ and $\mu_{v}(X,U)=\Pr(Y=2|X,U,V=v)$ The first estimator represents the typical case that the covariate $U$ is unmeasured, and the second represents the hypothetical case where $U$ is available, allowing for conditioning on a sufficient adjustment set. We consider both the typical case that the covariate $U$ is unmeasured, and the hypothetical case where $U$ is available, allowing for conditioning on a sufficient adjustment set. Finally, we include estimates from and $\widehat{\Psi}_{om}$ based on $\Psi_{om}$ from Proposition 1 and discussed further in section C that uses the full sample instead of restricting to those tested. Finally, in scenario 9, we demonstrate the double robustness property of $\widehat{\Psi}_{dr}^{*}$ by adjusting the data generation process for $V$ to be
@@
TND, om $\widehat{\Psi}^{*}_{om}$ Proposed estimator in expression 9. Nuisance terms $\mu_{v}^{*}(X)$ estimated via logistic regression of $Y^{*}$ on $(1,V,X,VX)^{\prime}$.
TND, ipw $\widehat{\Psi}^{*}_{ipw}$ Proposed estimator in expression 10. Nuisance term $\pi_{0}^{*}(X)$ estimated via logistic regression of $V$ on $(1,X)^{\prime}$ among those with $Y^{*}=0$.
TND, dr $\widehat{\Psi}^{*}_{dr}$ Proposed estimator in expression 11. Nuisance terms $\mu_{v}^{*}(X)$ and $\pi_{0}^{*}(X)$ estimated as above.
@@
cohort, U unmeasured $\widehat{\Psi}_{cohort}$ Standardization estimator in expression A13 assuming $U$ is unavailable. Nuisance terms $\mu_{v}(X)$ estimated via logistic regression of $Y$ on $(1,V,X,VX)^{\prime}$ in the full sample.
cohort, U measured $\widehat{\Psi}_{cohort,U}$ Standardization estimator in expression A14 assuming $U$ is available. Nuisance terms $\mu_{v}(X,U)$ estimated via logistic regression of $Y$ on $(1,V,X,VX,U)^{\prime}$ in the full sample.
@@
Table A4: Robustness of estimators to misspecification of nuisance functions: 1000 simulations of sample size $n=$15,000 from modified processes based on scenario 9 in Table 3. Bias is estimated mean bias, SE is estimated monte carlo standard error, Coverage is estimated coverage probability of 95% CI

### v3

@@ Appendix B Proofs of main identifiability results
    No effect of vaccination on testing negative and symptomatic among the vaccinated. That is, $\Pr(Y^{0}=1|V=1,X)=\Pr(Y^{1}=1|V=1,X).$
@@ Appendix D Doubly robust estimator under cohort sampling
$E\left[\Pr[Y=2|V=0,X]\dfrac{\Pr[Y=1|V=1,X]}{\Pr[Y=1|V=0,X]}\Big{|}V=1\right]$
@@
$E\left[\dfrac{(1-V)}{\Pr(V=1)}\mathbbm{1}(Y=2)\dfrac{\pi(1,X)}{1-\pi(1,X)}\right]$
@@ Appendix E Doubly robust estimator under TND sampling
Under TND sampling, i.e. $O_{TND}=\{(X_{i},V_{i},S_{i}=1,Y_{i}):i=1,\ldots,n\},$ with selection $S=\mathbbm{1}(Y\neq 0)$, we show in the main text that Equations 6 and 7, and by extension Equations A5 and A6, are still identifiable by
@@ G.2 Immortal time among vaccinated
Consider the emulation of a trial where events occurring in the first 28 days are excluded in both the vaccine arm and the control arm, as was specified in the original trials of the SARS-CoV-2 vaccines. This is often done because, biologically, it is believed that during this window the vaccine has not had time to provoke a sufficient immunological response and therefore provides no protection, although discarding these events may produce selection bias due to depletion of susceptibles when there is an effect. Regardless, attempts to emulate this result using a TND are challenged by the fact that the TND does not have a well-defined start of follow up and only the timing of vaccination and testing are known. Therefore, investigators instead discard cases where vaccination occurred less than 28 days prior to receiving a test, i.e. only among the vaccinated. This results in a form of immortal time bias as the vaccinated cannot have the outcome during this 28 day period. Assume for a second, to simplify exposition, that conditional on $X$ there is no other source of confounding. Note that under our framework, this would imply that
@@ Appendix H Identification when test is imperfect
In the main text, we assumed the availability of a perfect test such that $\widetilde{Y}_{i}=Y_{i}$ for all individuals where $\widetilde{Y}$ is the test result and $Y$ is the true value. In the absence of such a test, here we describe the potential for bias due to misclassification of case status which has also been described previously [sullivan_theoretical_2016]. Because most studies employ real-time reverse-transcription polymerase chain reaction (RT-PCR), false positives are unlikely as most RT-PCR tests have specificity approaching 100%. Therefore, misclassification of test-positives is probably rare, perhaps due to sample contamination or data entry errors. False negatives may be more likely as test sensitivity is generally lower due to the fact that (1) some of those infected with test-positive illness may not shed detectable virus or viral RNA; (2) some may seek care after shedding has ceased; or (3) sample quality may be compromised due to swab quality or inadequate storage. If sample collection is sufficient such that the source of the test-negative infection can be identified via RT-PCR, then sensitivity with respect to the test-positive illness can be improved by limiting to test-negative controls with an identified cause.
@@
    No effect of vaccination on test result That is, for any $v$ and $v^{\prime}$, we have
@@ Appendix I Simulation details
where $\mu_{v}(X)=\Pr(Y=2|X,V=v)$ and $\mu_{v}(X,U)=\Pr(Y=2|X,U,V=v)$ The first estimator represents the typical case that the covariate $U$ is unmeasured, and the second represents the hypothetical case where $U$ is available, allowing for conditioning on a sufficient adjustment set. We consider both the typical case that the covariate $U$ is unmeasured, and the hypothetical case where $U$ is available, allowing for conditioning on a sufficient adjustment set. Finally, we include estimates from and $\widehat{\Psi}_{om}$ based on $\Psi_{om}$ from Proposition 1 and discussed further in section C that uses the full sample instead of restricting to those tested. Finally, in scenario 8, we demonstrate the double robustness property of $\widehat{\Psi}_{dr}^{*}$ by adjusting the data generation process for $V$ to be
@@
TND, om $\widehat{\Psi}^{*}_{om}$ Proposed estimator in expression 10. Nuisance terms $\mu_{v}^{*}(X)$ estimated via logistic regression of $Y^{*}$ on $(1,V,X,VX)^{\prime}$.
TND, ipw $\widehat{\Psi}^{*}_{ipw}$ Proposed estimator in expression 11. Nuisance term $\pi_{0}^{*}(X)$ estimated via logistic regression of $V$ on $(1,X)^{\prime}$ among those with $Y^{*}=0$.
TND, dr $\widehat{\Psi}^{*}_{dr}$ Proposed estimator in expression 12. Nuisance terms $\mu_{v}^{*}(X)$ and $\pi_{0}^{*}(X)$ estimated as above.
@@
cohort, U unmeasured $\widehat{\Psi}_{cohort}$ Standardization estimator in expression A11 assuming $U$ is unavailable. Nuisance terms $\mu_{v}(X)$ estimated via logistic regression of $Y$ on $(1,V,X,VX)^{\prime}$ in the full sample.
cohort, U measured $\widehat{\Psi}_{cohort,U}$ Standardization estimator in expression A12 assuming $U$ is available. Nuisance terms $\mu_{v}(X,U)$ estimated via logistic regression of $Y$ on $(1,V,X,VX,U)^{\prime}$ in the full sample.
@@
Table A4: Robustness of estimators to misspecification of nuisance functions: 1000 simulations of sample size $n=$15,000 from modified processes based on scenario 8 in Table 3. Bias is estimated mean bias, SE is estimated monte carlo standard error, Coverage is estimated coverage probability of 95% CI